I have the quadratic 6x(-2x + 4) = 0. I get 2 solutions 0 or 2 but my vertex, b^2/2a = -24, this seems impossible. What solution do you get?
Found 2 solutions by rothauserc, Alan3354:
Answer by rothauserc(4718)

6x(-2x + 4) = 0
-12x^2 + 24x = 0
divide both sides of = by -12
x^2 - 2x = 0
factor
x(x - 2) = 0
x = 0 or 2
x coordinate for vertex is -(-2) / 2 = 1
y coordinate for vertex is 1 - 2 = -1
vertex is (1, -1)

Answer by Alan3354(69443)

I have the quadratic 6x(-2x + 4) = 0. I get 2 solutions 0 or 2 but my vertex, b^2/2a = -24, this seems impossible. What solution do you get?
-------------
x of the vertex is (0+2)/2 = 1
---
f(1) = 6*(-2 + 4) = 12
Vertex at (1,12)